Two cars start directly toward each other at the same time from towns which are 280 miles apart. They meet after 4 hours. one car travels 40 mph faster that the other car. Find the speed of each car
Answer by Alan3354(69443)

Two cars start directly toward each other at the same time from towns which are 280 miles apart. They meet after 4 hours. one car travels 40 mph faster that the other car. Find the speed of each car
--------------
280/4 = 70
---
x + y = 70
x - y = 40